A system absorbs 35J of heat and in the process it does 11J of work. (a) If the initial internal energy is 205J what is the final internal energy? (b) The system follows a different thermodynamic path to the same final state and does 15J of work, what is the heat transferred?

1
Expert's answer
2021-08-16T08:36:58-0400

a)


"U_f=205+35-11=229\ J"

b)


"Q=229-205-11=13\ J"